# SuperEvent: Cross-Modal Learning of Event-based Keypoint Detection for SLAM

### v1

@@ 4.4 Stereo Event-Visual Inertial SLAM
Table 3: Results on the small scale mocap sequences of TUM-VIE dataset, ATE in cm. The first three gray approaches require scale alignment with the ground truth, all other approaches estimate the absolute scale. Baseline numbers taken from [53, 26].
@@
ICRA’24 [54] Stereo E + IMU 3.85 18.90 failed 8.99 9.47 -
@@
We demonstrate the usefulness of SuperEvent with a downstream experiment. To date, frame-based visual odometry and SLAM systems largely outperform their event-based counterparts attributed to their decades-long lead in development. By integrating SuperEvent into the frame-based VI-SLAM system OKVIS2 [39], we enable this framework to process the keypoints detected in the event stream. We compare the performance of the resulting event-only system to other recent event-based E(I) Odometry and SLAM approaches on the TUM-VIE dataset [34] in Table 3. Moreover, note that the next best results of the data-driven Mono E(I)O systems DEVO and DEIO are achieved only after aligning the scale of the estimated trajectories with ground truth. Real-world applications, e.g., in VR and robotics, however, typically require accurate metric scale.
Figure 5: OKVIS2 + SuperEvent’s estimated trajectories of TUM-VIE sequence loop-floor0 with and without loop closure.
OKVIS2’s loop closure capability relies on long-term descriptor matching. While its effect on the small scale sequences is limited, we evaluate OKVIS2 + SuperEvent on the loop-floor sequences of the TUM-VIE dataset, since these are explicitly designed to test a system’s loop-closure capabilities. Thanks to SuperEvent, the OKVIS2 loop-closure feature works also reliably with events, as shown in Table 4. Figure 5 qualitatively compares the estimated trajectories of the sequence loop-floor0. Without loop closure, there is significant drift – albeit not out of the ordinary when compared to frame-based VI-SLAM. However, with loop closure enabled, the starting room is recognized and the drift is corrected. This demonstrates the robustness of SuperEvent, achieving successful descriptor matching even with minutes of time between observations.
Table 4: Effect of loop closure of OKVIS2 [39] + SuperEvent on TUM-VIE loop-floor sequences, ATE in cm.
loop-floor 0 1 2 3
[…]
Loop closure 4.96 4.64 8.92 4.74
W/o loop closure 132.11 161.92 116.00 129.17
@@ 5 Conclusion
We presented SuperEvent, a novel keypoint detector and descriptor for event streams. During its training, we consider the temporal dependence of keypoint appearance and therefore employ pseudo-labels detected on frames. Combined with the flexible input representation MCTS and a transformer backbone, it achieves SOTA performance on keypoint-based pose estimation. This further enables its integration into frame-based downstream applications. Therefore, SuperEvent unlocks event vision for existing frame-based systems relying on keypoints.
As future work, SuperEvent can be combined with a frame-based detector to realize systems that simultaneously process frames and events, exploiting their complementary advantages. Also, with a similar approach to SuperEvent, an event-based line detector and descriptor could be realized to improve the performance of downstream applications in man-made environments.
@@ A.1 Training Data
Table 5: Pose estimation after training SuperEvent with temporal matching data, homographic adaption data, and samples from using both methods.
@@
Homographic adaption + temporal matching 18.5 28.1 37.1 13.1 22.0 33.0
@@
We compare our proposed temporal matching approach with the state-of-the-art method homographic adaption introduced for frame-based keypoint detection and description [15] in Table 5. SuperEvent trained with temporal matching data exclusively outperforms models trained with homographic adaption data. Also, combining both approaches results does not improve model performance.
@@ A.2 Input Representation
[…]
@@ Appendix B Examples of Network Predictions and Pseudo-labels
Figure 7: Examples of predictions and pseudo-labels of data that was not used for training. Top (orange): Pseudo-labels generated by SuperPoint [15] + SuperGlue [60]; bottom (green): Predictions of SuperEvent.
Matched keypoints from SuperEvent on unseen sequences are shown in Figure 7. These sequences have been filtered out from the training data due to frame quality, low-light conditions or strong motion blur. The examples show the generalization ability of SuperEvent.
Appendix C Pose Estimation Experiment
@@
C.1 Benchmark Design
Pose Estimation requires reliable keypoint detection and matching and is therefore a common baseline for frame-based keypoint detectors [60, 65, 63, 21]. It also indicates the approaches’ usability for downstream applications such as Visual Odometry, Simultaneous Localization and Matching (SLAM), and Structure-from-Motion (SfM) that usually rely on keypoint-based pose estimation. Frame-based approaches are evaluated on datasets such as ScanNet [12] and MegaDepth [41] containing various images of the same scene with the associated ground truth camera poses allowing for a straightforward pose estimation evaluation. However, event datasets are usually temporally continuous sequences because of the asynchronous nature of the sensor. This raises the question, at which timestamps the pose estimation is evaluated. We define the evaluation benchmark as follows:
@@ C.1 Benchmark Design
[…]
@@
[…]
@@
[…]
@@
    For these selected prediction samples with associated ground truth measurements, we match the keypoint descriptors. Based on the resulting keypoint pairs and considering the lens distortion (unless the approach to evaluate already required a rectified event representation as input), the camera pose is estimated. We calculate the rotation difference and its angle of the axis-angle representation.
@@
We evaluate SuperEvent on the following sequences omitting the ones without sufficient rotation changes as well as calibration sequences.
@@
C.2 Why Not Homography Estimation?
@@
As long as there are always sufficiently many keypoints detected, this evaluation procedure might still be sensible for approaches that rely on basic nearest neighbor matching in pixel-space because some wrong matches do not lead to large errors. However, for approaches that rely on descriptor matching [47, 9, 10], only a few outliers with large errors have a serious negative impact on the reported score. Most downstream applications therefore employ outlier filtering, such as Random Sample Consensus (RANSAC)  [16]. Also, the approaches relying on descriptor matching [31, 18] employ RANSAC as their homography estimation benchmark. The RANSAC algorithm rejects outliers with a reprojection error greater than a pre-defined threshold $c_{\text{R}}$. Thus, this threshold becomes an upper bound to the estimated homography reprojection error since all keypoints with larger errors are filtered out. Thereby, the MRE score can be arbitrarily reduced by choosing a smaller $c_{\text{RE}}$, making it inappropriate for performance benchmarking.
This general problem does not only apply to RANSAC but to all approaches relying on some form of outlier removal as a post-processing step. Outliers, of course, are never completely avoidable, and it is a common procedure to remove them in downstream applications. Therefore, we decided to reproduce the frame-based benchmark of estimating the camera pose change in datasets with ground truth camera pose measurements – producing meaningful results for approaches with outlier removal.
Appendix D Stereo Event-Visual Intertial Odometry Experiment
Lastly, we visualize 2D projections of the trajectories estimated by SuperEvent integrated into OKVIS2 [39] yielding the reported results.
D.1 Small-after Scale Sequences
Figure 8 shows the trajectories from SuperEvent + OKVIS2 on the TUM-VIE [34] mocap-sequences. Since OKVIS2 is non-deterministic, we process every sequence 5 times and select the trajectory with median error.
@@ D.1 Small-after Scale Sequences
Figure 8: OKSVIS2 + SuperEvent’s trajectories on the TUM-VIE mocap sequences with ground truth trajectory (dashed) and absolute position error between the 3D trajectories shown by the color.
D.2 Large-Scale Sequences
The effect of loop closure on the trajectory estimation of OKVIS2 + SuperEvent can be seen in Figure 9. The loop closure is reliably detected on all 4 loop-floor sequences of the TUM-VIE dataset.

### v2

@@ 4.3 Stereo Event-Visual Inertial SLAM
Table 4: Results on the small scale mocap sequences of TUM-VIE dataset, ATE in cm. The first three gray results require scale alignment with the ground truth; all other approaches estimate the absolute scale. Baseline numbers taken from [54, 27].
@@
ICRA’24 [55] Stereo E + IMU 3.85 18.90 failed 8.99 9.47
@@
Table 5: Results on VECtor [18] large scale sequences. Our results are obtained with online calibration. Baseline numbers are taken from [27] and [36]. DEVO [36] cannot recover the absolute scale. Sequences marked ’–’ are not evaluated in [27]. ATE in cm.
corr.- corr.- units- units- school- school-
Method dolly walk dolly scooter dolly scooter
Ours 33.13 133.16 122.61 59.05 69.96 39.07
DEIO [27] 492.65 325.00 826.38 304.14
ESVO [68] failed failed failed failed 1371.0 983.0
We demonstrate the usefulness of SuperEvent with a downstream experiment. To date, frame-based visual odometry and SLAM systems largely outperform their event-based counterparts, attributed to their decades-long lead in development. By integrating SuperEvent into the frame-based VI-SLAM system OKVIS2 [40], we enable this framework to process the keypoints detected in the event stream. We compare the performance of the resulting event-only system to other recent event-based Odometry and SLAM approaches on the TUM-VIE dataset [35] in Table 4. Moreover, note that the next best results of the data-driven Mono E(I)O systems DEVO and DEIO are achieved only after aligning the scale of the estimated trajectories with ground truth. However, real-world applications, e.g., VR and robotics, typically require accurate metric scale. Results on the VECtor dataset [18] are listed in Table 5.
OKVIS2’s loop closure capability relies on long-term descriptor matching. While its effect on the small-scale sequences is limited, we evaluate OKVIS2 + SuperEvent on the loop-floor sequences of the TUM-VIE dataset, since these are explicitly designed to test a system’s loop-closure capabilities. Thanks to SuperEvent, the OKVIS2 loop-closure feature works reliably with events, as shown in Table LABEL:loop_floor. Figure LABEL:loop_floor0 qualitatively compares the estimated trajectories of the sequence loop-floor0. Without loop closure, there is significant drift – albeit not out of the ordinary compared to frame-based VI-SLAM. However, with loop closure enabled, the starting room is recognized and the drift is corrected. This demonstrates the robustness of SuperEvent, achieving successful descriptor matching even with minutes of time between observations.
@@ 5 Conclusion
We presented SuperEvent, a novel keypoint detector and descriptor for event streams. During its training, we consider the time-dependent keypoint appearance and employ pseudo-labels detected on frames. Combined with the flexible MCTS representation and a transformer backbone, it achieves SOTA performance on keypoint-based pose estimation. SuperEvent’s integration into frame-based downstream applications enables event vision in existing keypoint-based systems.
As future work, SuperEvent can be combined with a frame-based detector to realize systems that simultaneously process frames and events, exploiting their complementary advantages. Also, with a similar approach to SuperEvent, an event-based line detector could improve the performance of downstream applications in man-made environments.
@@ A.1 Training Data
Table 6: Pose estimation after training SuperEvent with temporal matching data, homographic adaptation data, and samples from both methods.
@@
Homographic adaptation + temp. matching 18.5 28.1 37.1 13.1 22.0 33.0
@@
We compare our proposed temporal matching approach with the state-of-the-art method homographic adaptation introduced for frame-based keypoint detection and description [15] in Table 6. SuperEvent trained with temporal matching data exclusively outperforms models trained with homographic adaptation data. Also, combining both approaches results does not improve model performance.
@@ A.2 Input Representation
Next, we compare our MCTS representation to time surfaces [38] and its variant Tencode [32] in Table 7. As shown in [32], with $\Delta t=$0.01\text{\,}\mathrm{s}$\$, the Tencode model achieves superior performance to the one using time surfaces. However, since Tencode is also used as a single channel tensor, an MCTS with a single time window size (but two channels) strictly separates polarities, thereby providing the model with additional information improving the performance further. Finally, more time windows increase the model’s robustness to fast or slow scene motions. Therefore, the 10-channel $\text{MCTS}_{5}$ with $\Delta t_{\{1,\ldots,N\}}=\{$0.001\text{\,}\mathrm{s}$,$0.003\text{\,}\mathrm{s}$,$0.01\text{\,}\mathrm{s}$,$0.03\text{\,}\mathrm{s}$,$0.1\text{\,}\mathrm{s}$\$\} enables the model to outperform all other variants.
@@ Appendix B Examples of Network Predictions and Pseudo-labels
Figure 6: Examples of predictions and pseudo-labels of data not used for training. Top (orange): pseudo-labels generated by SuperPoint [15] + SuperGlue [60]; bottom (green): predictions of SuperEvent.
Matched keypoints from SuperEvent on unseen sequences are shown in Figure 6. These sequences are held out during training, showing the generalization ability of SuperEvent.
Appendix C Event-based versus Frame-based Keypoint Matching
Since SuperEvent is not explicitly trained in scenarios where the quality of frame cameras degrades, such as fast scene motion and high dynamic range (HDR), we demonstrate its generalization ability by comparing SuperEvent’s event-based keypoint correspondences to related frame-based results.
C.1 Qualitative Comparison for Fast Motion and HDR
Figure 7 shows matches of the same scenes (not used for training) generated by SuperPoint [15] + SuperGlue [60] on the frames and by SuperEvent and brute-force matching on the event stream. While the quality of the frames under fast scene motion and HDR degrades, the event stream suffers less under these conditions, resulting in better and more equally distributed keypoint matches when using SuperEvent.
(a) HDR: DAVIS Driving Dataset 2020 [31]: rec1501614399
(b) HDR: Vision for Visibility Dataset [39]: Indoor Varying Robust
(c) Fast motion: UZH-FPV Drone Racing Dataset [14]: Outdoor 45° Downward Facing 1
(d) Fast motion: UZH-FPV Drone Racing Dataset [14]: Indoor 45° Downward Facing 14
Figure 7: Frame-based SuperPoint [15] + SuperGlue [60] (top), event-based SuperEvent (ours, bottom) on unseen sequences, showing the superior matching capabilities of SuperEvent by leveraging the higher quality of event data for scenes with HDR or fast scene motion.
C.2 Comparison in SLAM Downstream Task
We quantitatively compare the SLAM results from plain frame-based OKVIS2 [40] and our modified version after replacing the frame-based BRISK [41] detector with SuperEvent. Since event data degrades less than frames in such conditions, SuperEvent’s predictions improve OKVIS2’s estimations. Choosing a higher processing rate further boosts the performance: the visual overlap between MCTSs rises, and the effects of motion dependence are mitigated.
Table 9: Results on TUM-VIE [35] sequences with fast motion (mocap-shake) and low light (floor2-dark). ATE and RPE in cm; RPE for consecutive frames. Results marked are not representative due to discontinuous ground truth.
OKVIS2 [40] + SuperEvent (ours) frame-based
20 Hz 40 Hz 20 Hz
Sequence ATE RPE ATE RPE ATE RPE
mocap-shake 43.71 0.55 29.14 0.26 50.83 1.03
mocap-shake2 43.75 0.80 27.37 0.46 66.29 1.39
floor2-dark 9.58 2.51 9.37 1.23 failed failed
Appendix D Pose Estimation Experiment
@@
D.1 Benchmark Design
Pose Estimation requires reliable keypoint detection and matching and is therefore a common baseline for frame-based keypoint detectors [60, 65, 63, 22]. It also indicates the approaches’ usability for downstream applications such as Visual Odometry, Simultaneous Localization and Mapping (SLAM), and Structure-from-Motion (SfM) that usually rely on keypoint-based pose estimation. Frame-based approaches are evaluated on datasets such as ScanNet [12] and MegaDepth [42] containing various images of the same scene with the associated ground truth camera poses, allowing for a straightforward pose estimation evaluation. However, event datasets are usually temporally continuous sequences because of the sensor’s asynchronous nature. This raises the question, at which timestamps the pose estimation is evaluated. We define the evaluation benchmark as follows:
@@ D.1 Benchmark Design
In this experiment, we choose $c_{{\Delta\text{r},\text{max}}}=$45\text{\,}\text{\textdegree}$\$ $c_{{\Delta\tau}}=$2\text{\,}\mathrm{s}$\$ and $n_{\Delta\text{r}}=45$ to test various levels of difficulty while reducing the amount of samples without visual overlap. Our practical implementation executes the following steps for each sequence:
@@
    For all timestamps with associated ground truth pose measurements (usually between 100 and $200\text{\,}\mathrm{H}\mathrm{z}$), we generate keypoint and descriptor predictions. For tracking approaches, we assign a track ID as a scalar descriptor. Matching the same track ID reproduces the tracking result.
@@
    Next, we iteratively check for each ground truth sample if any of the subsequent ground truth samples within $c_{{\Delta\tau}}=$2\text{\,}\mathrm{s}$\$ yields an rotation change of at least $c_{{\Delta\text{r},\text{max}}}=$45\text{\,}\text{\textdegree}$\$\}. If this condition cannot be fulfilled, we skip the respective sample.
@@
    We match the keypoint descriptors for these selected prediction samples with associated ground truth measurements. The camera pose is estimated based on the resulting keypoint pairs and considering the lens distortion (unless the approach to evaluate already required a rectified event representation as input). We calculate the rotation difference and its angle of the axis-angle representation.
@@
We evaluate SuperEvent on the following sequences, omitting the ones without sufficient rotation changes as well as calibration sequences.
@@
D.2 Why Not Homography Estimation?
Figure 8: Homography estimation on HVGA ATIS [48] matching 400 keypoints per sample, $\Delta t=$50\text{\,}\mathrm{m}\mathrm{s}$\$, and RANSAC reprojection error thresholds of {0.1, 0.3, 1, 3, 10} pixels.
@@ D.2 Why Not Homography Estimation?
As long as there are always sufficiently many keypoints detected, this evaluation procedure might still be sensible for approaches that rely on basic nearest neighbor matching in pixel-space because some wrong matches do not lead to large errors. However, for approaches that rely on descriptor matching [48, 9, 10], only a few outliers with large errors have a serious negative impact on the reported score. Most downstream applications therefore employ outlier filtering, such as Random Sample Consensus (RANSAC)  [16]. Also, the approaches relying on descriptor matching [32, 19] employ RANSAC as their homography estimation benchmark. The RANSAC algorithm rejects outliers with a reprojection error greater than a pre-defined threshold $c_{\text{R}}$. Thus, this threshold is an upper bound to the estimated homography reprojection error since all keypoints with larger errors are filtered out. Thereby, the MRE score can be arbitrarily reduced by choosing a smaller $c_{\text{RE}}$, making it inappropriate for performance benchmarking.
This general problem applies not only to RANSAC but to all approaches relying on some form of outlier removal as a post-processing step. Outliers, of course, are never completely avoidable, and it is a common procedure to remove them in downstream applications. Therefore, we decided to reproduce the frame-based benchmark of estimating the camera pose change in datasets with ground truth camera pose measurements – producing meaningful results for approaches with outlier removal.
We illustrate this issue in Figure 8 where we plot the reprojection error after homography estimation for different RANSAC reprojection error thresholds over the ratio of matches classified as inliers: SuperEvent achieves similar reprojection errors as EventPoint [32] with RANSAC leading to far fewer outliers.
Appendix E Stereo Event-Visual Intertial SLAM Experiment
Table 10: RPE scores in cm for consecutive samples at 20 Hz of OKVIS2 [40] + SuperEvent on the TUM-VIE mocap sequences.
1d-trans 3d-trans 3dof desk desk2
0.07 0.14 0.10 1.39 1.25
Lastly, we visualize 2D projections of the trajectories estimated by SuperEvent integrated into OKVIS2 [40] yielding the reported results. In addition to the ATE reported in the main paper, we also report the RPE results of our method for completeness.
E.1 TUM-VIE Small-scale Sequences
Figure 9 shows the trajectories from SuperEvent + OKVIS2 on the TUM-VIE [35] mocap-sequences. Since OKVIS2 is non-deterministic, we process every sequence 5 times and select the trajectory with median error. We report the RPE scores in Table 10.
@@ E.1 TUM-VIE Small-scale Sequences
Figure 9: OKSVIS2 + SuperEvent’s trajectories on the TUM-VIE mocap sequences with ground truth trajectory (dashed) and absolute position error of the 3D trajectories shown by the color.
E.2 TUM-VIE Large-scale Sequences
The effect of loop closure on the trajectory estimation of OKVIS2 + SuperEvent can be seen in Figure 10. The loop closure is reliably detected on all 4 loop-floor sequences of the TUM-VIE dataset. Since the ground truth is not continuous, we do not report RPE scores for these sequences, as they lack interpretive value and are not comparable.
@@ E.2 TUM-VIE Large-scale Sequences
E.3 VECtor Large-scale Sequences
Figure 11 shows the trajectories from SuperEvent + OKVIS2 on the VECtor [18] large-scale sequences. We report RPE scores in Table 11.
Table 11: RPE scores in cm for consecutive samples at 20 Hz of OKVIS2 [40] + SuperEvent on the VECtor large-scale sequences.
corr.- corr.- units- units- school- school-
dolly walk dolly scooter dolly scooter
16.72 14.50 23.10 42.40 21.66 45.78
(a) corridors-dolly
(b) units-dolly
(c) school-dolly
(d) corridors-walk
(e) units-scooter
(f) school-scooter
Figure 11: OKSVIS2 + SuperEvent’s trajectories on the VECtor large-scale sequences with ground truth trajectory (dashed) and absolute position error of the 3D trajectories shown by the color.